# Entropy-based random models for hypergraphs

### v2

@@ Abstract
a sum whose generic addendum is 1 in just two cases: either $I_{i\alpha}=1$ and $I_{i\beta}=0$ or $I_{i\alpha}=0$ and $I_{i\beta}=1$. Since $d_{\alpha\beta}\sim\text{Bin}(N,2p(1-p))$, one finds that $P(d_{\alpha\beta}=0)=[1-2p(1-p)]^{N}$ and $\langle d_{\alpha\beta}\rangle=2p(1-p)N$, $\forall\>\alpha\neq\beta$. Notice that
$\lim_{T\to NL}P(d_{\alpha\beta}=0)$ $=\lim_{T\to 0}P(d_{\alpha\beta}=0)=1$ (15)
and
$\lim_{T\to NL}\langle d_{\alpha\beta}\rangle$ $=\lim_{T\to 0}\langle d_{\alpha\beta}\rangle=0$ (16)
as $(2T/NL)(1-T/NL)\to 0$ in both regimes: in words, the probability of observing parallel hyperedges progressively rises both as a consequence of having many 1s and as a consequence of having few 1s. Analogously for the expected Hamming distance.
Let us, now, evaluate the expected Hamming distance between any, two hyperedges $\alpha$ and $\beta$, within the asymptotic framework defined by letting $N\to+\infty$ while posing $T=O(L)$, i.e. $p=O(1/N)$. Since $p=T/NL$ and $h\equiv T/L$ remains finite, one finds that
@@
upon imposing $2h\geq 1$, i.e. that the expected Hamming distance between any, two hyperedges $\alpha$ and $\beta$ is at least 1, one derives what may be called resolution threshold, corresponding to $h_{r}^{\text{RHM}}=1/2$. In words, a value $p>p_{r}^{\text{RHM}}\equiv h_{r}^{\text{RHM}}/N=1/2N$ ensures that, on average, any, two hyperedges $\alpha$ and $\beta$ differ by at least one element (see also Appendix A).
III.1.3 3. Estimation of the overlap between hyperedges
Finally, let us move to considering the issue of overlapping hyperedges. To this aim, let us require that the probability for any, two hyperedges to overlap amounts to 1. In formulas,
$1-(1-p^{2})^{N}=1,$ (18)
a relationship leading to the equation
$(1-p^{2})^{N}\simeq 1-Np^{2}=0$ (19)
that holds true in the sparse regime and further implies
$h^{2}=N;$ (20)
such a value, firstly derived in Barthelemy (2022), individuates the so-called percolation threshold, corresponding to $h_{p}^{\text{RHM}}=\sqrt{N}$. In words, a value $p>p_{p}^{\text{RHM}}\equiv h_{p}^{\text{RHM}}/N=1/\sqrt{N}$ ensures that any, two hyperedges in our random hypergraph share at least one node.
Since $h_{r}^{\text{RHM}}\leq h_{f}^{\text{RHM}}\leq h_{p}^{\text{RHM}}$, such a result contributes to describe the following situation: by progressively rising the parameter $p$, three, different thresholds are met, i.e. i) the resolution threshold $p_{r}^{\text{RHM}}=1/2N$, ensuring that, on average, any, two hyperedges differ by at least one element; ii) the filling threshold $p_{f}^{\text{RHM}}=\ln L/N$, ensuring that the expected number of empty hyperedges is strictly less than one; iii) the percolation threshold $p_{p}^{\text{RHM}}=1/\sqrt{N}$, ensuring that any, two hyperedges overlap.
III.2 B. Heterogeneous benchmarks:
@@
The number of constraints can be enlarged to include the degrees, i.e. the sequence $\{k_{i}\}_{i=1}^{N}$, and the hyperdegrees, i.e. the sequence $\{h_{\alpha}\}_{\alpha=1}^{L}$. Counting the number of configurations on which both sequences match their empirical values is a hard task - although numerical recipes to shuffle the entries of a rectangular matrix, while preserving its marginals, exist Strona et al. (2014); Carstens (2015); Chodrow (2020). Solving the corresponding problem in the canonical framework is, instead, straightforward. Indeed, Shannon entropy maximisation leads to
$P(\mathbf{I})$ $=\frac{e^{-\sum_{i=1}^{N}\alpha_{i}k_{i}(\mathbf{I})-\sum_{\alpha=1}^{L}\beta_{\alpha}h_{\alpha}(\mathbf{I})}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{N}\alpha_{i}k_{i}(\mathbf{I})-\sum_{\alpha=1}^{L}\beta_{\alpha}h_{\alpha}(\mathbf{I})}}$
$=\frac{e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}(\alpha_{i}+\beta_{\alpha})I_{i\alpha}}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}(\alpha_{i}+\beta_{\alpha})I_{i\alpha}}}$
@@
with $e^{-\alpha_{i}}\equiv x_{i}$, $\forall\>i$, $e^{-\beta_{\alpha}}\equiv y_{\alpha}$, $\forall\>\alpha$ and $p_{i\alpha}\equiv x_{i}y_{\alpha}/(1+x_{i}y_{\alpha})$, $\forall\>i,\alpha$. According to such a model, the entries of the incidence matrix of a hypergraph are independent random variables that obey different Bernoulli distributions, i.e. $I_{i\alpha}\sim\text{Ber}(p_{i\alpha})$, $\forall\>i,\alpha$. As a consequence, both degrees and hyperdegrees obey Poisson-Binomial distributions, i.e. $k_{i}\sim\text{PoissBin}(L,\{p_{i\alpha}\}_{\alpha=1}^{L})$, $\forall\>i$ and $h_{\alpha}\sim\text{PoissBin}(N,\{p_{i\alpha}\}_{i=1}^{N})$, $\forall\>\alpha$ Saracco et al. (2017).
@@
The HCM reduces to a ‘partial’ Configuration Model Saracco et al. (2017) when either the degree or the hyperdegree sequence is left unconstrained (see also Appendix B). The canonical ensemble of each randomisation model (Table 2 in Appendix B sums up the set of constraints defining them) can be explicitly sampled by considering each entry of $\mathbf{I}$, drawing a real number $u_{i\alpha}\in U[0,1]$ and posing $I_{i\alpha}=1$ if $u_{i\alpha}\leq p_{i\alpha}$, $\forall\>i,\alpha$.
The HCM is formally equivalent to the Bipartite Configuration Model (BiCM) Saracco et al. (2015). Such an identification is guaranteed by our focus on non-simple hypergraphs.
III.2.1 1. Estimation of the number of empty hyperedges
@@
$p_{0}^{\alpha}\equiv\prod_{i=1}^{N}(1-p_{i\alpha}).$ (26)
Such a quantity leads to the probability of observing at least one, empty hyperedge, now reading
$p_{1}\equiv 1-\prod_{\alpha=1}^{L}\left[1-p_{0}^{\alpha}\right]=1-\prod_{\alpha=1}^{L}\left[1-\prod_{i=1}^{N}(1-p_{i\alpha})\right],$ (27)
and to the expected number of empty hyperedges
$\langle N_{0}\rangle\equiv\sum_{\alpha=1}^{L}p_{0}^{\alpha}=\sum_{\alpha=1}^{L}\prod_{i=1}^{N}(1-p_{i\alpha}).$ (28)
As in the previous section, let us, now, inspect the behaviour of the aforementioned quantities on the ensemble induced by the HCM, as the density of 1s in the incidence matrix varies. Although it depends on (the heterogeneity of) the sets of coefficients $\{x_{i}\}_{i=1}^{N}$ and $\{y_{\alpha}\}_{\alpha=1}^{L}$, general conclusions can still be drawn within a simpler framework. To this aim, let us consider the functional form reading
@@
where the vector of fitnesses $\{f_{i}\}_{i=1}^{N}$ accounts for the heterogeneity of nodes, the vector of fitnesses $\{g_{\alpha}\}_{\alpha=1}^{L}$ accounts for the heterogeneity of hyperedges and $z$ tunes the density of 1s in the incidence matrix [^1] Partial Configuration Models are recovered upon posing either $f_{i}=1$, $\forall\>i$ or $g_{\alpha}=1$, $\forall\>\alpha$.. Within such a framework, the fitnesses of the nodes and the fitnesses of the hyperedges can be drawn from any distribution. The dense and sparse regimes are, now, defined by the positions $z\to+\infty$ and $z\to 0$, respectively. In the dense case, one finds
$\lim_{z\to+\infty}p_{0}^{\alpha}$ $=\lim_{z\to+\infty}\prod_{i=1}^{N}(1-p_{i\alpha})$
$=\lim_{z\to+\infty}\prod_{i=1}^{N}\left(1-\frac{zf_{i}g_{\alpha}}{1+zf_{i}g_{\alpha}}\right)=0,$ (30)
@@
a relationship inducing $p_{1}\xrightarrow[]{z\to+\infty}0$ and $\langle N_{0}\rangle\xrightarrow[]{z\to+\infty}0$: in words, the probability of observing empty hyperedges progressively vanishes as the density of 1s increases. Consistently, in the sparse case one finds
$\lim_{z\to 0}p_{0}^{\alpha}$ $=\lim_{z\to 0}\prod_{i=1}^{N}(1-p_{i\alpha})$
$=\lim_{z\to 0}\prod_{i=1}^{N}\left(1-\frac{zf_{i}g_{\alpha}}{1+zf_{i}g_{\alpha}}\right)=1,$ (31)
@@
a relationship inducing $p_{1}\xrightarrow[]{z\to 0}1$ and $\langle N_{0}\rangle\xrightarrow[]{z\to 0}L$. In words, the probability of observing empty hyperedges progressively rises as the density of 1s decreases.
III.2.2 2. Estimation of the number of parallel hyperedges
@@
Let us, now, focus on the issue of parallel hyperedges. As in the case of the RHM, we consider the Hamming distance between the columns representing the two hyperedges $\alpha$ and $\beta$. Since, now, $d_{\alpha\beta}\sim\text{PoissBin}(N,\{q_{i}^{\alpha\beta}\}_{i=1}^{N})$, where $q_{i}^{\alpha\beta}\equiv p_{i\alpha}(1-p_{i\beta})+p_{i\beta}(1-p_{i\alpha})$ with $p_{i\alpha}=zf_{i}g_{\alpha}/(1+zf_{i}g_{\alpha})$ and $p_{i\beta}=zf_{i}g_{\beta}/(1+zf_{i}g_{\beta})$, one finds that $P(d_{\alpha\beta}=0)=\prod_{i=1}^{N}(1-q_{i}^{\alpha\beta})$ and $\langle d_{\alpha\beta}\rangle=\sum_{i=1}^{N}q_{i}^{\alpha\beta}$, $\forall\>\alpha\neq\beta$. Notice that
$\lim_{z\to+\infty}P(d_{\alpha\beta}=0)$ $=\lim_{z\to 0}P(d_{\alpha\beta}=0)=1$ (32)
@@
$\lim_{z\to 0}\langle d_{\alpha\beta}\rangle$ $=\lim_{z\to 0}\langle d_{\alpha\beta}\rangle=0$ (33)
as $q_{i}^{\alpha\beta}\to 0$ in both regimes. As in the case of the RHM, the probability of observing parallel hyperedges progressively rises both as a consequence of having many 1s and as a consequence of having few 1s. Analogously for the expected Hamming distance.
When considering heterogeneous benchmarks, closed-form expressions for the asymptotic estimates can be provided only in specific cases (see also Appendix B).
Figure 2: Left panel: trend of $p_{1}\xrightarrow[]{N\to+\infty}1-(1-e^{-h})^{L}$, i.e. the probability of observing at least one empty hyperedge as a function of $T/NL$; evaluating it in correspondence of $h_{f}^{\text{RHM}}=\ln L\simeq 6.907$ (whose related, critical density value $p_{f}^{\text{RHM}}=h_{f}^{\text{RHM}}/N=\ln L/N\simeq 0.023$ individuates a steep transition separating the sparse regime from the dense one) returns the value $0.63231-(1-1/L)^{L}=$ 0.6323. Right panel: trend of $p_{1}$ as defined by Eq. (27). Notice that the trend induced by the HCM is shifted on the right with respect to the trend induced by the RHM: this is probably due to the presence of small fitnesses increasing the probability of observing at least one, empty hyperedge, hence requiring a larger value of $z$ to let $p_{1}$ vanish. The dense (sparse) regime is recovered for large (small) values of $T/NL$. Each dot represents an average taken over an ensemble of $1000$ configurations (explicitly sampled from either the RHM or the HCM) and is accompanied by the corresponding $95\%$ confidence interval.
IV IV. Results
IV.1 A. Hypergraphs in the dense and sparse regime

### v3

@@ Methods
a sum whose generic addendum is 1 in just two cases: either $I_{i\alpha}=1$ and $I_{i\beta}=0$ or $I_{i\alpha}=0$ and $I_{i\beta}=1$. Since $d_{\alpha\beta}\sim\text{Bin}(N,2p(1-p))$, one finds that
$p_{\mathbin{/\!/}}^{\alpha\beta}\equiv P(d_{\alpha\beta}=0)=[1-2p(1-p)]^{N}$ (22)
and that
$\langle d_{\alpha\beta}\rangle=2p(1-p)N$ (23)
$\forall\>\alpha\neq\beta$. Since $p(1-p)=(T/NL)(1-T/NL)$, one finds that
$\lim_{T\to NL}p(1-p)$ $=\lim_{T\to 0}p(1-p)=0,$ (24)
i.e. $p(1-p)$ vanishes in both regimes, a result further implying that both $P(d_{\alpha\beta}=0)\xrightarrow[]{T\to NL}1$ and $P(d_{\alpha\beta}=0)\xrightarrow[]{T\to 0}1$ and that both $\langle d_{\alpha\beta}\rangle\xrightarrow[]{T\to NL}0$ and $\langle d_{\alpha\beta}\rangle\xrightarrow[]{T\to 0}0$: in words, the probability of observing parallel hyperedges progressively rises both as a consequence of having many $1$s and as a consequence of having few $1$s. Analogously for the expected Hamming distance.
Let us, now, evaluate the expected Hamming distance between any two hyperedges $\alpha$ and $\beta$ within the asymptotic framework defined by letting $N\to+\infty$ while posing $T=O(L)$ equivalently, defined by posing $p=O(1/N)$. Since $p=T/NL$ and $h\equiv T/L$ remains finite, one finds that
@@
upon imposing $2h\geq 1$, i.e. that the expected Hamming distance between any two hyperedges $\alpha$ and $\beta$ is at least 1, one derives what may be called resolution threshold, corresponding to $h_{r}^{\text{RHM}}\equiv 1/2$. In words, a value $p>p_{r}^{\text{RHM}}=h_{r}^{\text{RHM}}/N=1/2N$ ensures that, on average, any, two hyperedges $\alpha$ and $\beta$ differ by at least one element.
To derive a global condition on the total number of parallel hyperedges, note that, although the overlaps between pairs of hyperedges cannot be treated as i.i.d. random variables, the expected number of parallel hyperedges can still be computed explicitly. Upon posing $p_{\mathbin{/\!/}}\equiv p_{\mathbin{/\!/}}^{\alpha\beta}$, it reads
${\color[rgb]{0,0,0}\langle N_{\mathbin{/\!/}}\rangle=\sum_{\alpha=1}^{L}\sum_{\begin{subarray}{c}\beta=1\\ \beta>\alpha\end{subarray}}^{L}p_{\mathbin{/\!/}}=\frac{L(L-1)}{2}p_{\mathbin{/\!/}}.}$ (26)
Considering that $p_{\mathbin{/\!/}}\xrightarrow[]{N\to+\infty}e^{-2h}$ and imposing $\langle N_{\mathbin{/\!/}}\rangle\leq 1$, i.e. that the expected number of parallel hyperedges is at most 1, one derives what may be called a multiple resolution threshold, corresponding to $h_{m}^{\text{RHM}}\equiv\ln L-\ln\sqrt{2}\lesssim h_{f}^{\text{RHM}}$. In words, a value
${\color[rgb]{0,0,0}p>p_{m}^{\text{RHM}}=\frac{h_{m}^{\text{RHM}}}{N}=\frac{\ln L}{N}-\frac{\ln\sqrt{2}}{N}\lesssim\frac{\ln L}{N}=\frac{h_{f}^{\text{RHM}}}{N}=p_{f}^{\text{RHM}}}$ (27)
(also) ensures that the expected number of parallel hyperedges in our random hypergraph is strictly less than one.
Estimation of the percolation threshold
The two thresholds derived in the previous subsections emerge in consequence of the attempts of solving the problems related to the appearance of empty as well as parallel hyperedges. Remarkably, a third threshold exists: known as percolation threshold, it was, first, derived in [12], following the definition according to which any two hyperedges are said to be connected if they share at least one node. Here, we re-derive the percolation threshold by considering the ‘hypergraph to graph’ projection (see also below): more formally, the total number of nodes shared by hyperedge $\alpha$ with any other hyperedge reads
${\color[rgb]{0,0,0}\sigma_{\alpha}=\sum_{i=1}^{N}\sum_{\begin{subarray}{c}\beta=1\\ \beta\neq\alpha\end{subarray}}^{L}I_{i\alpha}I_{i\beta},}$ (28)
its expected value being
${\color[rgb]{0,0,0}\langle\sigma_{\alpha}\rangle=N(L-1)p^{2}\simeq NLp^{2};}$ (29)
imposing $\langle\sigma_{\alpha}\rangle=1$ leads to find the value
${\color[rgb]{0,0,0}p_{p}^{\text{RHM}}=\frac{h_{p}^{\text{RHM}}}{N}=\frac{1}{\sqrt{NL}}}$ (30)
which, in turn, induces the value $h_{p}^{\text{RHM}}\equiv\sqrt{N/L}$. In words, a value $p>p_{p}^{\text{RHM}}$ ensures that any two hyperedges in our random hypergraph share, on average, at least one node.
Heterogeneous benchmarks:
@@
Microcanonical formulation
The number of constraints can be enlarged to include the degrees, i.e. the sequence $\{k_{i}\}_{i=1}^{N}$, and the hyperdegrees, i.e. the sequence $\{h_{\alpha}\}_{\alpha=1}^{L}$. Although counting the number of configurations on which both sequences match their empirical values is a hard task, numerical recipes that shuffle the entries of a rectangular matrix, while preserving its marginals, exist [36, 37, 9, 38]. It should be, however, noticed that, if not carefully implemented, algorithms of the kind may lead to a non-uniform exploration of the space of configurations [39, 40]; moreover, the issue concerning the time needed to collect a sufficiently large number of configurations should be addressed even in presence of an ergodic system. A recent proposal is that of extending the traditional Curveball algorithm to hypergraphs [38].
Canonical formulation
Solving the corresponding problem in the canonical framework is, instead, straightforward. Indeed, Shannon entropy maximisation leads to
$P(\mathbf{I})$ $=\frac{e^{-\sum_{i=1}^{N}\alpha_{i}k_{i}(\mathbf{I})-\sum_{\alpha=1}^{L}\beta_{\alpha}h_{\alpha}(\mathbf{I})}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{N}\alpha_{i}k_{i}(\mathbf{I})-\sum_{\alpha=1}^{L}\beta_{\alpha}h_{\alpha}(\mathbf{I})}}=\frac{e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}(\alpha_{i}+\beta_{\alpha})I_{i\alpha}}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}(\alpha_{i}+\beta_{\alpha})I_{i\alpha}}}$
@@
with $e^{-\alpha_{i}}\equiv x_{i}$, $\forall\>i$, $e^{-\beta_{\alpha}}\equiv y_{\alpha}$, $\forall\>\alpha$ and $p_{i\alpha}\equiv x_{i}y_{\alpha}/(1+x_{i}y_{\alpha})$, $\forall\>i,\alpha$. According to such a model, the entries of the incidence matrix of a hypergraph are independent random variables that obey different Bernoulli distributions, i.e. $I_{i\alpha}\sim\text{Ber}(p_{i\alpha})$, $\forall\>i,\alpha$. As a consequence, both degrees and hyperdegrees obey Poisson-Binomial distributions, i.e. $k_{i}\sim\text{PoissBin}(L,\{p_{i\alpha}\}_{\alpha=1}^{L})$, $\forall\>i$ and $h_{\alpha}\sim\text{PoissBin}(N,\{p_{i\alpha}\}_{i=1}^{N})$, $\forall\>\alpha$ [24].
Parameter estimation
@@
The HCM reduces to a ‘partial’ Configuration Model [24] when either the degree or the hyperdegree sequence is left unconstrained (see also Section A in the Appendix). The canonical ensemble of each randomisation model (Table 1 in the Appendix sums up the sets of constraints defining them) can be explicitly sampled by considering each entry of $\mathbf{I}$, drawing a real number $u_{i\alpha}\in U[0,1]$ and posing $I_{i\alpha}=1$ if $u_{i\alpha}\leq p_{i\alpha}$, $\forall\>i,\alpha$.
The HCM is formally equivalent to the Bipartite Configuration Model (BiCM) [35]. Such an identification is guaranteed by our focus on non-simple hypergraphs.
Estimation of the number of empty hyperedges
@@
$p_{\emptyset}^{\alpha}\equiv\prod_{i=1}^{N}(1-p_{i\alpha})$ (36)
while the expected number of empty hyperedges, now, reads
$\langle N_{\emptyset}\rangle\equiv\sum_{\alpha=1}^{L}p_{\emptyset}^{\alpha}=\sum_{\alpha=1}^{L}\prod_{i=1}^{N}(1-p_{i\alpha}).$ (37)
As previously done, let us inspect the behaviour of the aforementioned quantities on the ensemble induced by the HCM as the density of $1$s in the incidence matrix varies. Although it depends on (the heterogeneity of) the sets of coefficients $\{x_{i}\}_{i=1}^{N}$ and $\{y_{\alpha}\}_{\alpha=1}^{L}$, general conclusions can be still drawn within a simpler framework. To this aim, let us consider the functional form reading
@@
where the vector of fitnesses $\{f_{i}\}_{i=1}^{N}$ accounts for the heterogeneity of nodes, the vector of fitnesses $\{g_{\alpha}\}_{\alpha=1}^{L}$ accounts for the heterogeneity of hyperedges and $z$ tunes the density of $1$s in the incidence matrix - ‘partial’ Configuration Models are recovered upon posing either $f_{i}=1$, $\forall\>i$ or $g_{\alpha}=1$, $\forall\>\alpha$.) Within such a framework, the fitnesses of the nodes and the fitnesses of the hyperedges can be drawn from any distribution. The dense and sparse regimes are, now, defined by the positions $z\to+\infty$ and $z\to 0$, respectively. In the dense case, one finds
$\lim_{z\to+\infty}p_{\emptyset}^{\alpha}=\lim_{z\to+\infty}\prod_{i=1}^{N}\left(1-\frac{zf_{i}g_{\alpha}}{1+zf_{i}g_{\alpha}}\right)=0,$ (39)
@@
a relationship inducing $\langle N_{\emptyset}\rangle\xrightarrow[]{z\to+\infty}0$: in words, the probability of observing empty hyperedges progressively vanishes as the density of $1$s increases. Consistently, in the sparse case one finds
$\lim_{z\to 0}p_{\emptyset}^{\alpha}=\lim_{z\to 0}\prod_{i=1}^{N}\left(1-\frac{zf_{i}g_{\alpha}}{1+zf_{i}g_{\alpha}}\right)=1,$ (40)
@@
a relationship inducing $\langle N_{\emptyset}\rangle\xrightarrow[]{z\to 0}L$: in words, the probability of observing empty hyperedges progressively rises as the density of $1$s decreases.
For what concerns the filling threshold, a derivation that is similar-in-spirit to the one carried out for the case of the RHM can be sketched. Let us pose ourselves in the sparse regime: since $1-p_{i\alpha}\simeq e^{-p_{i\alpha}}$, the probability for the generic hyperedge to be empty satisfies the chain of relationships
${\color[rgb]{0,0,0}p_{\emptyset}^{\alpha}=\prod_{i=1}^{N}(1-p_{i\alpha})\simeq\prod_{i=1}^{N}e^{-p_{i\alpha}}=e^{-\sum_{i=1}^{N}p_{i\alpha}}=e^{-h_{\alpha}};}$ (41)
consistently, the expected number of empty hyperedges becomes
${\color[rgb]{0,0,0}\langle N_{\emptyset}\rangle=\sum_{\alpha=1}^{L}p_{\emptyset}^{\alpha}\simeq\sum_{\alpha=1}^{L}e^{-h_{\alpha}}}$ (42)
and imposing $\langle N_{\emptyset}\rangle\leq 1$, i.e. that the expected number of empty hyperedges is at most 1, one derives a global condition to be satisfied by the hyperdegrees. In general terms, the aforementioned condition leads to require $e^{-h_{\alpha}}=O(1/L)$, i.e. $h_{\alpha}=O(\ln L)$, $\forall\>\alpha$ and $p_{i\alpha}=O(\ln L/N)$, $\forall\>i,\alpha$.
A comparison with simple graphs: estimating the number of isolated nodes
Coming to traditional graphs, the aim is, now, to estimate the number of isolated nodes. In this case, $1-p_{ij}\simeq e^{-p_{ij}}$ and the probability for the generic node $i$ to be isolated reads
${\color[rgb]{0,0,0}q_{0}^{i}=\prod_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}(1-p_{ij})\simeq\prod_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}e^{-p_{ij}}=e^{-\sum_{j(\neq i)=1}^{N}p_{ij}}=e^{-k_{i}};}$ (43)
consistently, the expected number of isolated nodes becomes
${\color[rgb]{0,0,0}\langle N_{0}\rangle=\sum_{i=1}^{N}q_{0}^{i}\simeq\sum_{i=1}^{N}e^{-k_{i}}}$ (44)
and imposing $\langle N_{0}\rangle\leq 1$, i.e. that the expected number of isolated nodes is at most 1, one derives a global condition to be satisfied by the degrees. In general terms, the aforementioned condition leads to require $e^{-k_{i}}=O(1/N)$, i.e. $k_{i}=O(\ln N)$, $\forall\>i$ and $p_{ij}=O(\ln N/N)$, $\forall\>i<j$.
Estimation of the number of parallel hyperedges
@@
As in the case of the RHM, we consider the Hamming distance between the columns representing the two hyperedges $\alpha$ and $\beta$. Since, now, $d_{\alpha\beta}\sim\text{PoissBin}(N,\{q_{i}^{\alpha\beta}\}_{i=1}^{N})$, where $q_{i}^{\alpha\beta}\equiv p_{i\alpha}(1-p_{i\beta})+p_{i\beta}(1-p_{i\alpha})$ with $p_{i\alpha}=zf_{i}g_{\alpha}/(1+zf_{i}g_{\alpha})$ and $p_{i\beta}=zf_{i}g_{\beta}/(1+zf_{i}g_{\beta})$, one finds that
$P(d_{\alpha\beta}=0)=\prod_{i=1}^{N}(1-q_{i}^{\alpha\beta})$ (45)
and that
$\langle d_{\alpha\beta}\rangle=\sum_{i=1}^{N}q_{i}^{\alpha\beta}$ (46)
$\forall\>\alpha\neq\beta$. Since
$\lim_{z\to+\infty}q_{i}^{\alpha\beta}$ $=\lim_{z\to 0}q_{i}^{\alpha\beta}=0,$ (47)
i.e. $q_{i}^{\alpha\beta}$ vanishes in both regimes, one finds that both $P(d_{\alpha\beta}=0)\xrightarrow[]{z\to+\infty}1$ and $P(d_{\alpha\beta}=0)\xrightarrow[]{z\to 0}1$ and that both $\langle d_{\alpha\beta}\rangle\xrightarrow[]{z\to+\infty}0$ and $\langle d_{\alpha\beta}\rangle\xrightarrow[]{z\to 0}0$: as in the case of the RHM, the probability of observing parallel hyperedges progressively rises both as a consequence of having many $1$s and as a consequence of having few $1$s. Analogously for the expected Hamming distance.
For what concerns the resolution threshold, a derivation that is similar-in-spirit to the one carried out for the case of the RHM can be sketched. Let us pose ourselves in the sparse regime and consider that $1-q_{i}^{\alpha\beta}\simeq e^{-q_{i}^{\alpha\beta}}\simeq e^{-(p_{i\alpha}+p_{i\beta})}$. As a consequence
${\color[rgb]{0,0,0}p_{\mathbin{/\!/}}^{\alpha\beta}\equiv P(d_{\alpha\beta}=0)=\prod_{i=1}^{N}(1-q_{i}^{\alpha\beta})\simeq\prod_{i=1}^{N}e^{-q_{i}^{\alpha\beta}}\simeq e^{-\sum_{i=1}^{N}(p_{i\alpha}+p_{i\beta})}=e^{-(h_{\alpha}+h_{\beta})}}$ (48)
@@
${\color[rgb]{0,0,0}\langle d_{\alpha\beta}\rangle=\sum_{i=1}^{N}q_{i}^{\alpha\beta}\simeq\sum_{i=1}^{N}(p_{i\alpha}+p_{i\beta})=h_{\alpha}+h_{\beta}.}$ (49)
Imposing $\langle d_{\alpha\beta}\rangle\geq 1$, i.e. that the expected Hamming distance between any two hyperedges $\alpha$ and $\beta$ is at least 1, amounts to require that $P(d_{\alpha\beta}=0)\leq e^{-1}$ thus recovering the same condition holding true in the case of the RHM where, in fact, $p_{\mathbin{/\!/}}\equiv p_{\mathbin{/\!/}}^{\alpha\beta}\xrightarrow[]{N\to+\infty}e^{-2h}$.
The expected number of parallel hyperedges, now, reads
@@
${\color[rgb]{0,0,0}\langle N_{\mathbin{/\!/}}\rangle=\sum_{\alpha=1}^{L}\sum_{\begin{subarray}{c}\beta=1\\ \beta>\alpha\end{subarray}}^{L}p_{\mathbin{/\!/}}^{\alpha\beta}\simeq\sum_{\alpha=1}^{L}\sum_{\begin{subarray}{c}\beta=1\\ \beta>\alpha\end{subarray}}^{L}e^{-(h_{\alpha}+h_{\beta})};}$ (50)
upon imposing $\langle N_{\mathbin{/\!/}}\rangle\leq 1$, i.e. that the expected number of parallel hyperedges is at most 1, one derives a global condition to be satisfied by the hyperdegrees. In general terms, the aforementioned condition leads to require $e^{-(h_{\alpha}+h_{\beta})}=O(1/L^{2})$, i.e. $h_{\alpha}=O(\ln L)$, $\forall\>\alpha$ and $p_{i\alpha}=O(\ln L/N)$, $\forall\>i,\alpha$.
Estimation of the percolation threshold
For what concerns the percolation threshold, the expected value of the total number of nodes shared by hyperedge $\alpha$ with any other hyperedge, now, reads
${\color[rgb]{0,0,0}\langle\sigma_{\alpha}\rangle=\sum_{i=1}^{N}\sum_{\begin{subarray}{c}\beta=1\\ \beta\neq\alpha\end{subarray}}^{L}p_{i\alpha}p_{i\beta}}$ (51)
and imposing $\langle\sigma_{\alpha}\rangle=1$ leads to a global condition to be satisfied. In general terms, the aforementioned condition leads to require $p_{i\alpha}p_{i\beta}=O(1/NL)$, i.e. $p_{i\alpha}=O(1/\sqrt{NL})$, $\forall\>i,\alpha$.
Results
Hypergraphs in the dense and sparse regime